OpenStudy (kirbykirby):

Let \(f:\mathbb{C} \rightarrow \mathbb{C}\) be continuous and assume that \(f(z)=f(2z)\) for all \(z \in \mathbb{C}\). Show that \(f\) is constant.

OpenStudy (anonymous):

If you're looking for a proof here, I *think* you can try this by contradiction. Assume \(f(z)\) is non-constant, perhaps by saying it's monotonically increasing or decreasing for some interval. Is it possible then for \(f(z)=f(2z)\) to be true? Just an idea, I don't know if it'd take you anywhere.

OpenStudy (anonymous):

I should take that back. I'm not even sure monotonicity applies to complex-valued functions...

OpenStudy (kirbykirby):

Yeah I'm not too sure about that either

OpenStudy (kinggeorge):

If you had the fact that \(f\) was holomorphic, I might be able to do something, but I don't see anything (right now) if it's just continuous.

OpenStudy (kinggeorge):

I think I have an idea. Since it's continuous everywhere, it's continuous at 0. So let \(\{z_i\}_{i=1}^\infty\) be a sequence of complex numbers approaching 0 along a straight line so that \(z_i=2z_{i+1}\). So since \(f\) is continuous, \[\lim_{n\to\infty}f(z_n)=f(0).\]But from the way we defined our sequence, \(\lim_{n\to\infty}f(z_n)=f(z_1)\). So \(f(z_1)=f(0)\). Now, notice that we can choose \(z_1\) arbitrarily in \(\mathbb{C}\) and define the same sequence, and it will always approach \(0\). But that means that \(f(0)=f(z)\) for all \(z\in\mathbb{C}\), so the function must be constant.

OpenStudy (kirbykirby):

how do you say that \(\lim_{n\rightarrow \infty} f(z_n)=f(z_1)\)

OpenStudy (kinggeorge):

That's from the condition that \(f(z)=f(2z)\). The way we defined the sequence, we know that \(f(z_i)=f(z_j)\) for all \(i,j\in\mathbb{Z}^+\). So if we say \(f(z_1)=L\), then we could rewrite the limit as \(\lim_{n\to\infty}L\). But \(L\) is a constant, so the limit is clearly equal to \(L=f(z_1)=f(0)\).

OpenStudy (kirbykirby):

Gosh I feel really dumb but I'm not sure I quite understand :( Maybe because I have trouble understand how complex sequences work? Honestly our prof spent like 30 mins on that... can we imagine that they work similarly as in real-valued functions? Also, if you constructed this particular sequences, how does it imply it's true in all cases?

OpenStudy (kinggeorge):

Convergent sequences in the complex numbers are slightly different from sequence in the real numbers, but the particular sequence I constructed is almost exactly like a sequence in the real numbers since every point in the sequence lies on the same line. As for continuous functions, I'm using the definition: A function \(f\) is continuous at a point \(c\) if for any sequence \(\{x_1\}_{n=1}^\infty\) that converges to \(c\), then the sequence \(\{f(x_n)\}_{n=1}^\infty\) converges to \(f(c)\) (so the limit as \(n\to\infty\) of \(f(x_n)\) will be \(f(c)\). It's true in all cases, since if I choose ANY \(z_1\in\mathbb{C}\), then define the sequence \(z_{i+1}=z_i/2\), the sequence \(\{z_i\}_{i=1}^\infty\) will approach the origin. Since the function is continuous, we use the above definition to see that \(f(z_i)\) approaches \(f(0)\) as \(i\to\infty\). Finally, from the condition that \(f(z)=f(2z)\), or equivalently \(f(z)=f(z/2)\), we notice that \(f(z_i)\) is constant for all \(i\in\mathbb{Z}^+\). So the sequence must approach that same constant. But we already know it approaches \(f(0)\), so \(f(0)\) must be that constant. So \(f(0)=f(z_1)\) for all \(z_1\in\mathbb{C}\), so \(f\) must be constant.

OpenStudy (kirbykirby):

ohhh oh my I had to read this MANY times but it has finally sunk in! I get it now. Thank you sooooo much!!!

OpenStudy (kinggeorge):

You're welcome. Out of curiosity, what is this for? Complex analysis? Real analysis? Or something different?

OpenStudy (kirbykirby):

It is for complex analysis

OpenStudy (kinggeorge):

Nice. You'll be surprised at how much nicer things work in the complex plane rather than the real line.

OpenStudy (kirbykirby):

Hehe good to know